# Cosmic muon flux attenuation methods for superconducting qubit experiments

### v4

@@ Abstract
We propose and demonstrate two practical mitigation methods to attenuate the cosmic muon flux, compatible with experiments involving superconducting qubits: shallow underground sites and specific device orientation. Using a specifically-built cosmic muon detector, we identify underground sites, widely present in urban environments, where significant attenuation of cosmic muon flux, up to a factor 35 for 100-meter depths, can be attained. Furthermore, we employ two germanium wafers in an above-ground laboratory, each equipped with a particle sensor, to show how the orientation of the chip with respect to the sky affects the amount and type of energy deposited on the substrate by ionizing radiation. We observe that the horizontal detector sees more counts at lower energy, while the vertical one is impacted by more particles at higher energy. The methods here described proposed ways to directly understand and reduce the effects of cosmic rays on qubits by attenuating the source of this type of decoherence, complementing existing on-chip mitigation strategies. We expect that both on-chip and off-chip methods combined will become ubiquitous in quantum technologies based on superconducting qubit circuits.
@@ Introduction
Ionizing radiation events are currently not the dominant source of superconducting qubit decoherence [4]. Nevertheless, increasing evidence points to this radiation being responsible for qubit device stability and performance [1], through its effects lingering on the qubit substrate in the form of phonons [5], quasiparticles [6], and the scrambling of two-level defects [7]. Therefore, techniques to mitigate the impact of ionizing radiation events are already necessary today. While on-chip mitigation strategies, such as phonon [8] or quasiparticle traps [9, 10, 11], are being developed, it is also necessary to understand the physics behind ionizing radiation, particularly the cosmic muon events, in order to find ways to attenuate its flux. In this work, we propose two such methods. The first method is to consider shallow underground sites of moderate depths of up to 100 m, where significant attenuation factors of muon flux have been characterized [12, 13, 14, 15]. The second method studies device-orientation dependence with respect to the zenith angle, since muons reaching the Earth’s surface have a strong angular dependence due to their absorption through the atmosphere. Specifically, for the average energy of muons reaching the Earth’s surface (4 GeV) a cosine-square dependence is expected.
We perform a muon flux characterization at different depths using a home-built muon detector and we verify the data collected with tailored Geant4 simulations. We provide a detailed description of the detector so that other groups can develop their own.
In parallel, we measured the amount of energy deposition by ionizing radiation at cryogenic temperatures for two identical germanium wafers: this experiment shows how the orientation of the chip with respect to the sky can strongly change the energy deposition within the substrate, mainly due to the strong angular dependence of cosmic rays reaching the device.
@@
This work is organized as follows: in Sec. 1 we give an introduction of the physics of cosmic rays and its connection with qubits. In Sec. 2, we describe our muon detector and present its characterization. Sec. 3 presents our main results of the dependence of cosmic muon flux on the rock depth for shallow sites along with the Geant4 simulations developed to cross-check the data collected. Sec. 4 presents the difference in energy spectra recorded in two germanium wafers, one oriented parallel and one orthogonal with respect to the sky. Finally, we provide conclusions and future perspectives in Sec. 5.
@@ 1.1.1 Cosmic rays at sea level
Muons at sea level exhibit a rate of approximately 1 count per minute per cm2 of horizontal surface and a mean energy around 4 GeV [19].
@@ 1.1.2 Cosmic rays underground
In underground laboratories, the muon flux decreases almost exponentially with the thickness of the rock overburden above the laboratory [20] (see Sec. 3). Normally, to compare the depth of different laboratories, since the rock composition varies, the overburden is converted in meter water equivalent (m.w.e.) 111The conversion between m.w.e. and real meters depends on the rock composition, with typical values in the range of 2.4-3 m.w.e./m..
@@
These shallow laboratories offer various advantages: they reduce the amount of direct muon interactions with the sample, they reduce the number of fast neutrons generated by spallation events caused by cosmic protons and muons, and they also reduce the cosmogenic activation of materials constituting or surrounding the experimental setup [25, 26].
@@ 1.2 Qubit decoherence due to ionizing radiation
The second mechanism has been recently identified to be a scrambling effect on the two-level systems [7] (TLSs) believed to reside mostly at the superconductor-insulator interface of the qubit circuit [46]. These TSLs display a very broad frequency spectrum, often coming into resonance with qubits they couple to, leading to either fluctuations in the relaxation time () [7] or gaps in the energy spectrum [47]. The scrambling effect causes shifts in the frequency of TLSs and thus it is responsible for both frequency noise in qubits as well as fluctuations in and the decoherence time () [48].
@@
Both mechanisms are particularly harmful as they cause correlated noise [5] which cannot be accounted for in currently explored quantum error correcting methods such as the surface code [49]. Another possible mechanism not studied are direct hits on the superconducting material, similar to the detection mechanism of single photons in superconducting nanowire photon detectors [50]. The deleterious effect of cosmic rays was also demonstrated on superconducting resonators placed in the deep underground laboratory in Gran Sasso [2]. Other evidence of the impact of ionizing radiation has been demonstrated on multi-qubit devices, where correlated charge fluctuating events between neighboring qubits were observed [6], and in the stability of fluxons in superconducting loops [51].
@@
The first direct evidence of the impact of ionizing radiation on a superconducting qubit used a radioactive source installed directly inside of a dilution refrigerator mixing chamber stage [1]. A transmon qubit placed in the vicinity of the source experimented a drastic decrease of its lifetime due to the activity of the source. Based on those measurements, an upper bound of the transmon qubit lifetime due to background ionizing radiation of more than 3 milliseconds was established, with an estimated 40% of the external background events corresponding to cosmic rays. Other works using fluxonium qubits have reached millisecond-range lifetimes [52] and coherences [53], possibly limited by quasiparticles (QPs) partially generated by ionizing radiation as well as stray microwave fields [54].
@@
Assuming one is able to eliminate all background radioactivity by proper shielding inside and outside the cryostat, only cosmic rays will impact qubit coherence. A decrease of the cosmic muon flux would lead to a reduction of the steady-state amount of excess QPs, given the reduction of energetic phonons in the substrate. The relation between high energy impacts and QPs is complex and depends on many material properties. In addition, a reduction of cosmic muon flux would also reduce the rate of TLS fluctuating frequency, leading to more stable qubit frequencies and coherence times.
@@
Environmental radioactivity from neighboring sources inside the dilution refrigerator are progressively being identified [55]. Attention will be placed in future experimental setups and laboratory spaces towards more radiopure systems, such as those employed in complex, low-background experiments [23]. Eventually, the only significant remaining source of ionizing radiation will be the cosmic rays, which, as stated in the previous section, cannot be easily screened above ground. Therefore, mitigation strategies are needed to suppress the effects of cosmic rays. In this work, we propose two methods: the screening obtained in shallow underground sites (Sec. 3) and the specific device orientation (Sec. 4). Other initiatives in the field have been addressing the same problem with direct on-chip mitigation strategies, such as lower-gap superconducting quasiparticle traps [9], phonon traps [8], or suspending the device on a membrane [56]. All strategies have evidenced a decrease on the impact of ionizing events on the qubits and/or resonators employed as sensors. Very likely, the final solution will become a combination of both external as well as internal mitigation methods.
@@ 2 Compact muon detector
In this section, we present a compact and portable cosmic muon detector that we built to characterize the flux of muons at shallow underground sites of different depths.
@@ 2.2 Description of the detector
The detector has been built in a very compact setup for easy transportation and fits into a standard size “executive” suitcase (see Fig. 1). It is powered by a 12 V lead battery or alternatively connected to the 220 V AC electrical network. In this second case, the battery is recharged in parallel. From either the 12 V from the battery or the 220 V AC power supply delivering 12 V output, a power management circuitry delivers all the needed voltages (see Fig. 2).
@@ 2.3 Detector characterization
In our detector, typical values measured for the single rates at ground level after taking data for 20 hours are CPM. As mentioned before, s, so that the calculated accidental coincidence rate is
@@
For shallow and deep underground sites, this error becomes very relevant as it is comparable or larger than the actual reading (see Sec. 3).
@@ 3 Shallow Underground Sites
As explained in Sec. 1, experiments requiring the lowest possible noise background level locate their sensors in deep underground laboratories around the world, where the rock overburden screens most of the incoming cosmic radiation. Typical depths of underground laboratories are in the 1 km range, obtaining screening factors of cosmic muon flux of [58, 59]. Due to their remote location and scarcity, these underground stations do not represent a practical solution for improving the quality of any qubit type affected by cosmic rays. As the lifetime of superconducting qubits limited by ionizing radiation lies already in the millisecond range for transmons [1], it may not be necessary to attain the screening factors achieved in deep underground laboratories and, instead, much shallower sites [14] may already offer a significant enough screening to bring qubit quality to a level desirable for applications such as error correction and sensing.
@@ 3.1 Shallow Underground Survey
We have identified commonly-found locations in city environments providing significant screening factors for cosmic radiation. These are underground stations and side tunnels in underground highways. For the particular case of the city of Barcelona, we have conducted measurements in two such locations using our portable cosmic muon detector described in Sec. 2. The first site is the underground station ‘El Coll-La Teixonera’ located under a hill at a depth of approximately 100 meters (see Fig. 4).
@@
As a reference of a more realistic depth where qubit laboratories could be easily located, we also performed measurements in a much shallower site of 6 m deep in the sub-basement of the ALBA synchrotron located near the Bellaterra campus333https://www.cells.es/.
The results are shown in Fig. 6. All measurements were taken over 48 hours to minimize statistical uncertainty (smaller than the data markers on the plot). Above ground, we measure on average 41.38 CPM. The TMB and Vallvidrera shallow sites investigated display a muon count of 1.4 CPM and 1.1 CPM, respectively, achieving screening factors of 29 and 36 with respect to ground. The 6 m shallower site achieves a rate of 20 CPM, representing a screening factor of 2. In all data, we have already subtracted the spurious coincidence counts from the apparatus dead time, as detailed in Sec. 2.3.
Inside the Vallvidrera tunnels, we have collected data for single events in our detector with the result of 1047.4 CPM. This represents a factor 1.68 higher than above ground in a usual office space, which is consistent with the presence of heavy minerals in the tunnels generating a higher amount of radioactivity. This is also consistent with measurements in deep underground laboratories. Therefore, it is necessary to equip qubit measurement setups with screening against this type of radioactivity, such as a lead shield. Note that given the low amount of cosmic ray flux in this shallow site, the spallation events discussed in Sec. 1.1 would also be reduced.
Assuming one is able to screen environmental radioactivity using lead shielding and employ radiopure setups, as is usual in underground sites, the screening factor obtained at 100 m depths would lead to a decrease of events to a level of s-1cm-2. Using typical device dimensions for superconducting qubits of cm-2, a device placed in such a shallow site would experience an average of 0.018 counts per minute, or 0.3 mHz, compared to 0.54 counts per minute, or 9 mHz, above ground.
@@ 3.2 Data cross-check with Geant4 Simulations
In order to cross-check the data obtained, we performed Geant4 [60] simulations of the portable detector above ground and underground, for the two configurations (vertical and horizontal) described in Sec. 2.4.
@@
Due to geometric limitations, muons with zenith angles greater than 70 degrees are not simulated, which initially would lead to an expected deficit in the simulation compared to the data, with a reduction of approximately 35. However, an overabundance is observed in the simulated data. A plausible explanation is that muons at larger angles are strongly attenuated by the terrain’s topography.
@@ 4.1 Experimental Methods
To test the orientation effect of impacts from cosmic radiation, we analyze the data collected by two light detectors developed for the CROSS [62] and BINGO [63] experiments. The detectors were operated simultaneously in a dilution refrigerator located in the above-ground laboratory IJCLab in Orsay (France) for 16.28 hours. For the vertical orientation data, we refer to detector CuCo2, for the horizontal orientation detector GeSph6, see Fig. 7. The two detectors consist of a germanium wafer with identical (45450.3) mm3 dimensions. Each one is equipped with a Neutron-Transporation-Doped (NTD) germanium thermistor [64], a sensor that measures temperature variations induced by particle interactions within the wafer. The NTD ensures high stability and detection efficiency along with high-energy resolution and a low-energy threshold for particle detection. Each NTD is glued directly on top of the germanium wafers using a standardized procedure and they are wire-bonded using a 25m gold wire to copper-capton-copper pads thermally connected to the coldest stage of the dilution refrigerator. Hence, the gold wires provide the link between the thermal bath and the NTDs as well as the electrical connections to measure the temperature variations of the detector. In fact, the temperature of the NTD is read out by measuring its voltage drop while applying a constant bias current through the gold wires.
The energy calibration of the detectors was performed using a mix of features visible in the spectrum of each detector and shining LED light through a fiber optic.
Multiple methods exist to establish the energy threshold of bolometers [65]: we determine the threshold by starting from the distribution of baselines without a distinguishable signal (noise-only): we take 1,000 baselines, randomly selected within the collected data, extract for each an amplitude with the same methodology used for particle pulses, and plot these amplitudes. The distribution can be approximated to a Gaussian and can be accuratrly fit to extract its FWHM. The vertical detector has a FWHM baseline of 0.666 0.010 keV while the horizontal detector has a

### v6

@@ Abstract
We propose and demonstrate two practical mitigation methods to attenuate the cosmic muon flux, compatible with experiments involving superconducting qubits: shallow underground sites and device orientation. Using a specifically-built cosmic muon detector, we identify underground sites, widely present in urban environments, where significant attenuation of cosmic muon flux, up to a factor 35 for 100-meter depths, can be attained. Furthermore, we employ two germanium wafers in an above-ground laboratory, each equipped with a particle sensor, to show how the orientation of a chip with respect to the sky affects the amount and type of energy deposited on the substrate by ionizing radiation. We observe that the horizontal detector sees more counts at lower energy, while the vertical one is impacted by more particles at higher energy. The methods here described proposed ways to directly understand and reduce the effects of cosmic rays on qubits by attenuating the source of this type of decoherence, complementing existing on-chip mitigation strategies. We expect that both on-chip and off-chip methods combined will become ubiquitous in quantum technologies based on superconducting qubit circuits.
@@ Introduction
Ionizing radiation events are currently not the dominant source of superconducting qubit decoherence [4]. Nevertheless, increasing evidence points to this radiation being responsible for qubit device stability and performance [1], through its effects lingering on the qubit substrate in the form of phonons [5], quasiparticles [6], and the scrambling of two-level defects [7]. Therefore, techniques to mitigate the impact of ionizing radiation events are already necessary today. While on-chip mitigation strategies, such as phonon [8] or quasiparticle traps [9, 10, 11], are being developed, it is also necessary to understand the physics behind ionizing radiation, particularly the cosmic muon events, in order to find ways to attenuate its flux. In this work, we propose two such methods. The first method is to identify shallow underground sites at accessible areas such as urban environments, considering the attenuation factors of muon flux already characterized at such moderate depths [12, 13, 14, 15], which may already set no limitations to qubit coherence. The second method studies device-orientation dependence with respect to the zenith angle, since muons reaching the Earth’s surface have a strong angular dependence due to their absorption through the atmosphere. Specifically, for the average energy of muons reaching the Earth’s surface (4 GeV) a cosine-square dependence is expected.
We perform a muon flux characterization at different depths using a home-built muon detector and verify the data collected with tailored Geant4 simulations. We also provide a detailed description of the detector so that it can be reproduced.
In parallel, we measured the amount of energy deposition by ionizing radiation at cryogenic temperatures for two identical germanium wafers: this experiment showed how the orientation of the chip with respect to the sky can strongly change the energy deposition within the substrate, mainly due to the strong angular dependence of cosmic rays reaching the device.
@@
This work is organized as follows: in Section 1 we give an introduction of the physics of cosmic rays and its connection with qubits. In Section 2, we describe our specific muon detector and present its characterization. Section 3 presents our main results of the dependence of cosmic muon flux on the rock depth for shallow sites along with the Geant4 simulations developed to cross-check the data collected. Section 4 presents the difference in energy spectra recorded in two germanium wafers, one oriented parallel with respect to the sky (zenith angle of 90 degrees), the other orthogonal with respect to the sky (zenith angle of 0 degrees). Finally, we provide conclusions and future perspectives in Section 5.
@@ 1.1.1 Cosmic rays at sea level
Muons, which are the main focus of this work, are elementary particles of the lepton family with an electric charge of -1e, spin 1/2, and a 105.66 MeV/c2 mass [18]. They are also unstable, with a mean lifetime of 2.2 s [18]. At sea level, muons exhibit a rate of approximately 1 count per minute per cm2 of horizontal surface and a mean energy around 4 GeV [19].
@@ 1.1.2 Cosmic rays underground
In underground laboratories, the muon flux decreases almost exponentially with the thickness of the rock overburden above the laboratory [20] (see Section 3). Normally, to compare the depth of different laboratories, since the rock composition varies, the overburden is converted in meter water equivalent (m.w.e.) 111The conversion between m.w.e. and real meters depends on the rock composition, with typical values in the range of 2.4-3 m.w.e./m..
@@
These shallow laboratories offer various advantages: they reduce the amount of direct muon interactions with the sample, they reduce the number of fast neutrons generated by spallation events caused by cosmic protons and muons, and they also reduce the cosmogenic activation of materials constituting or surrounding the experimental setup [25, 26]. In addition, and contrary to deep underground sites, shallow laboratories present far less difficulties to be established at accessible areas such as urban environments.
@@ 1.2 Qubit decoherence due to ionizing radiation
The second mechanism has been recently identified to be a scrambling effect on the two-level systems [7] (TLSs), believed to reside mostly at the superconductor-insulator interface of the qubit circuit [46]. These TSLs display a very broad frequency spectrum, often coming into resonance with qubits they couple to, leading to either fluctuations in the relaxation time () [7] or gaps in the energy spectrum [47]. The scrambling effect causes shifts in the frequency of TLSs and thus it is responsible for both frequency noise in qubits as well as fluctuations in and the decoherence time () [48].
@@
Both mechanisms are particularly harmful as they cause correlated noise [5] which cannot be accounted for in currently explored quantum error correcting methods such as the surface code [49]. Another possible mechanism not currently studied are direct hits on the superconducting material, similar to the detection mechanism of single photons in superconducting nanowire photon detectors [50]. The deleterious effect of cosmic rays was also demonstrated on superconducting resonators placed in the deep underground laboratory in Gran Sasso [2]. Other evidence of the impact of ionizing radiation has been demonstrated on multi-qubit devices, where correlated charge fluctuating events between neighboring qubits were observed [6], and in the stability of fluxons (the magnetic flux quanta) in superconducting loops [51].
@@
The first direct evidence of the impact of ionizing radiation on a superconducting qubit used a radioactive source installed directly inside of a dilution refrigerator mixing chamber stage [1]. A transmon qubit placed in the vicinity of the source experimented a drastic decrease of its lifetime due to the activity of the source. Based on those measurements, an upper bound of the transmon qubit lifetime due to background ionizing radiation of more than 3 milliseconds was established, with an estimated 40% of the external background events corresponding to cosmic rays. Other works using fluxonium qubits have reached millisecond-range lifetimes [52] and coherences [53], possibly limited by quasiparticles (QPs) partially generated by ionizing radiation as well as by stray microwave fields [54].
@@
Assuming one is able to eliminate all background radioactivity by proper shielding inside and outside the cryostat, only cosmic rays will impact qubit coherence. A decrease of the cosmic muon flux would lead to a reduction of the steady-state amount of excess QPs, given the reduction of energetic phonons in the substrate. The relation between high energy impacts and QP density is complex and depends on several material properties. In addition, a reduction of cosmic muon flux would also reduce the rate of TLS fluctuating frequency, leading to more stable qubit frequencies and coherence times.
@@
Environmental radioactivity from neighboring sources inside the dilution refrigerator are progressively being identified [55]. Attention will be placed in future experimental setups and laboratory spaces towards more radiopure systems, such as those employed in complex, low-background experiments [23]. Eventually, the only significant remaining source of ionizing radiation will be the cosmic rays, which, as stated in the previous section, cannot be easily screened above ground. Therefore, mitigation strategies are needed to suppress the effects of cosmic rays. In this work, we propose two methods: the screening obtained in shallow underground sites (Section 3) and the specific device orientation (Section 4). Other initiatives in the field have been addressing the same problem with direct on-chip mitigation strategies, such as lower-gap superconducting quasiparticle traps [9], phonon traps [8], or suspending the device on a membrane [56]. All strategies have evidenced a decrease on the impact of ionizing events on the qubits and/or resonators employed as sensors. Very likely, the final solution will become a combination of both external as well as internal mitigation methods. Finally, a novel unique strategy proposes to use a cryogenic veto system compatible with the operation of superconducting qubit devices [57].
@@ 2 Compact muon detector
In this section, we present a compact and portable cosmic muon detector that we purposely built to characterize the flux of muons at shallow underground sites of different depths.
@@ 2.2 Description of the detector
The detector has been built in a very compact setup for easy transportation and fits into a standard size “executive” suitcase (see Fig. 1). It is powered by a 12 V lead battery or alternatively connected to the electrical network. In this second case, the battery is recharged in parallel. From either the 12 V from the battery or the AC power supply delivering 12 V output, a power management circuitry delivers all the needed voltages (see Fig. 2).
@@ 2.3 Detector characterization
In our detector, typical values measured for the single rates at ground level after taking data for 20 hours are Counts Per Minute (CPM). As mentioned before, s, so that the calculated accidental coincidence rate is
@@
For shallow and deep underground sites, this error becomes very relevant as it is comparable or larger than the actual reading (see Section 3).
@@ 3 Shallow Underground Sites
As explained in Section 1, experiments requiring the lowest possible noise background level locate their sensors in deep underground laboratories around the world, where the rock overburden screens most of the incoming cosmic radiation. Typical depths of underground laboratories are in the 1 km range, obtaining screening factors of cosmic muon flux of [59, 60]. Due to their remote location and scarcity, these underground stations do not represent a practical solution for improving the quality of any qubit type affected by cosmic rays. As the lifetime of superconducting qubits limited by ionizing radiation lies already in the millisecond range for transmons [1], it may not be necessary to attain the screening factors achieved in deep underground laboratories and, instead, much shallower sites [14] located in accessible areas may already offer a significant enough screening to bring qubit quality to a level desirable for applications such as error correction and sensing.
@@ 3.1 Shallow Underground Survey
We have identified commonly-found locations in city environments providing significant screening factors for cosmic radiation. These are underground stations and side tunnels in underground highways. For the particular case of the city of Barcelona, we have conducted measurements in two such locations using our portable cosmic muon detector described in Section 2. The first site is the underground station ‘El Coll-La Teixonera’ located under a hill at a depth of approximately 100 meters (see Fig. 4).
@@
As a reference of a more realistic depth where qubit laboratories could be easily located, we also performed measurements in a much shallower site 6 m deep in the sub-basement of the ALBA synchrotron located near the Bellaterra campus333https://www.cells.es/.
The results are shown in Fig. 6. All measurements were taken over 48 hours to minimize statistical uncertainty (smaller than the data markers on the plot). Above ground, we measure on average 41.38 CPM. The TMB and Vallvidrera shallow sites investigated display a muon count of 1.4 CPM and 1.1 CPM, respectively, achieving screening factors of 29 and 36 with respect the above ground measurements. The 6 m shallower site achieves a rate of 20 CPM, representing a screening factor of 2. In all data, we have already subtracted the spurious coincidence counts from the apparatus dead time, as detailed in Section 2.3. In Fig. 6 we show the counts per cm2 per minute, calculated from the CPM divided by the effective area of the scintillators in cm2.
Inside the Vallvidrera tunnels, we have collected data for single events in our detector with the result of 1047.4 CPM. This represents a factor 1.68 higher than above ground in a usual office space, which is consistent with the presence of heavy minerals in the tunnels generating a higher amount of radioactivity. This is also consistent with measurements in deep underground laboratories. Therefore, it is necessary to equip qubit measurement setups with screening against this type of radioactivity, such as a lead shield. Note that given the low amount of cosmic ray flux in this shallow site, the spallation events discussed in Section 1.1 would also be reduced.
Assuming one is able to screen environmental radioactivity using lead shielding and employ radiopure setups, as is usual in underground sites, the screening factor obtained at 100 m depths would lead to a decrease of events to a level of s-1cm-2. Using typical device dimensions for superconducting qubit chips of cm-2, a device placed in such a shallow site would experience an average of 0.018 counts per minute, or 0.3 mHz, compared to 0.54 counts per minute, or 9 mHz, above ground.
@@ 3.2 Data cross-check with Geant4 Simulations
In order to cross-check the data obtained, we performed Geant4 [61] simulations of the portable detector above ground and underground, for the two configurations (vertical and horizontal) described in Section 2.4.
@@
Due to geometric limitations, muons with zenith angles greater than 70 degrees are not simulated, which initially would lead to an expected deficit in the simulation compared to the data, with a reduction of approximately 35. However, an overabundance is observed in the simulated data. A plausible explanation is that muons at larger angles are strongly attenuated by the terrain’s topography, leading to a lower count by our detector compared to the simulated muons.
@@ 4.1 Experimental Methods
To test the orientation effect of impacts from cosmic radiation, we analyze the data collected by two light detectors developed for the CROSS [63] and BINGO [64] experiments. The detectors were operated simultaneously in a dilution refrigerator located in the above-ground laboratory IJCLab in Orsay (France), for 16.28 hours. For the vertical orientation data, we refer to detector CuCo2, while for the horizontal orientation we refer to detector GeSph6, see Fig. 7. The two detectors consist of a germanium wafer with identical (45450.3) mm3 dimensions. Each one is equipped with a Neutron-Transportation-Doped (NTD) germanium thermistor [65], a sensor that measures temperature variations induced by particle interactions within the wafer. The NTD ensures high stability and detection efficiency along with high-energy resolution and a low-energy threshold for particle detection. Each NTD is glued directly on top of the germanium wafers using a standardized procedure and they are wire-bonded using a 25 µm gold wire to copper-capton-copper pads thermally connected to the coldest stage of the dilution refrigerator. Hence, the gold wires provide the link between the thermal bath and the NTDs as well as the electrical connections to measure the temperature variations of the detector. In fact, the temperature of the NTD is read out by measuring its voltage drop while applying a constant bias current through the gold wires.
The energy calibration of the detectors was performed using a mix of features visible in the spectrum of each detector and shining a 600 nm LED placed at room temperature through an optical fiber to the Mixing Chamber of the dilution refrigerator. The LED is pulsed for 2 µs to obtain a characteristic line used to perform a rough energy calibration. This setup is well calibrated [66], and as such the peak visible in Fig. 8 is expected to fall at an energy of 100-300 keV. These fast LED pulses result in a single narrow peak in the energy spectrum, since this bolometer cannot temporally distinguish between an energy deposition caused by a single source or multiple particles at this time scale. As such, all the photons emitted by the LED and subsequently absorbed by the bolometer will appear as a single pulse in the detector.
Multiple methods exist to establish the energy threshold of bolometers [67]: we determine the threshold by starting from the distribution of baselines without a distinguishable signal (noise-only): we take 1,000 baselines, randomly selected within the collected data, extract for each an amplitude with the same methodology used for particle pulses, and plot these amplitudes. The distribution can be approximated to a Gaussian and can be accurately fit to extract its FWHM. The vertical detector has a FWHM baseline of 0.666 0.010 keV while the horizontal detector has a